# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 2.2. Generalization to $\Phi$-Sobolev
    $\frac{p(r)}{r}$ is a non-decreasing function of $r$;
  3. (3)
    $u(\delta):=e^{\delta}\Phi(e^{-\delta}),\;\delta\in\mathbb{R},$ is a convex function of $\delta$.
Theorem 6 can be easily generalized to the following $\Phi$-entropy version.
Theorem 7 (From Nonlinear $\Phi$-Sobolev to Nonlinear $\Phi$-Transport).
Let $\Phi\in\mathcal{DC}_{\infty}$. Assume that the $\theta$-nonlinear $\Phi$-Sobolev inequality
$\theta(D_{\Phi}(\pi\|\nu))\leq I_{\Phi}(\pi\|\nu),\;\forall\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
holds for some function $\theta:[0,\infty]\to[0,\infty]$. Further assume that the equation $\partial_{t}\rho=Lp(\rho)$ for an initial distribution $\mu\in P_{2}^{\mathrm{ac}}(\mathcal{M})$ admits a solution $\mu_{t}=\rho_{t}\nu$ satisfying Assumptions (a) to (f) in Theorem 24.2 of [65] (called “smooth” in [65]), where $L=\Delta-\nabla V\cdot\nabla$ is a Laplace-type operator. Then, it holds that
(2.10) $\mathsf{W}_{2}\left(\mu,\mu_{t}\right)\leq\int_{D_{\Phi}(\mu_{t}\|\nu)}^{D_{\Phi}(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
In particular, letting $t\to\infty$ yields that
(2.11) $\mathsf{W}_{2}\left(\mu,\nu\right)\leq\int_{0}^{D_{\Phi}(\mu\|\nu)}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
3. From Transport to Variant Isoperimetry
3.1. From Transport to Variant Isoperimetry
We now consider a general setting for the isoperimetric problem. Let $\mathcal{X}$ and $\mathcal{Y}$ be Polish spaces (i.e., separable completely metrizable spaces, including complete Riemannian manifolds). Let $c:\mathcal{X}\times\mathcal{Y}\to[0,+\infty)$ be lower semi-continuous, which is called a cost function. Denote $\mathcal{X}^{n}$ (resp. $\mathcal{Y}^{n}$) as the $n$-fold product space of $\mathcal{X}$ (resp. $\mathcal{Y}$). Given $c$, we consider an additive cost function $c_{n}$ on $\mathcal{X}^{n}\times\mathcal{Y}^{n}$ given by
$c_{n}(x,y):=\sum_{i=1}^{n}c(x_{i},y_{i})\quad\mbox{for }(x,y)\in\mathcal{X}^{n}\times\mathcal{Y}^{n}.$
For a set $A\subseteq\mathcal{X}^{n}$, denote its $t$-enlargement under $c_{n}$ as
(3.1) $A^{c_{n}\leq t}:=\{\mathbf{y}\in\mathcal{Y}^{n}:c_{n}(\mathbf{x},\mathbf{y})\leq t,\mathbf{x}\in A\}.$
In particular, if $c=d^{2}$ on the Riemannian manifold $\mathcal{M}$, then $c_{n}^{1/2}$ corresponds to the product distance $d_{n}$. Let $\nu_{X}$ and $\nu_{Y}$ be probability measures on $\mathcal{X}$ and $\mathcal{Y}$ respectively, and let $\nu_{X}^{\otimes n}$ and $\nu_{Y}^{\otimes n}$ be the corresponding product measures.
Define the isoperimetric function as for $a\in[0,1],t\geq 0$,
(3.2) $\Gamma_{n}(a,t):=\inf_{A:\nu_{X}^{\otimes n}(A)\geq a}\nu_{Y}^{\otimes n}(A^{c_{n}\leq t}).$
We set $a=e^{-n\alpha}$ and $t=n\tau$. That is, we focus on exponentially small sets and enlarge them by a linear cost. Define the isoperimetric exponent as for $\alpha,\tau\geq 0$,
(3.3) $E_{n}(\alpha,\tau)$ $:=-\frac{1}{n}\ln\Gamma_{n}(e^{-n\alpha},n\tau).$
The optimal transport cost between two probability measures $\pi_{X}$ and $\pi_{Y}$ is defined as
(3.4) $\mathsf{C}(\pi_{X},\pi_{Y}):=\min_{\pi\in\Pi(\pi_{X},\pi_{Y})}\int c(x,y){\text{\rm d}}\pi(x,y),$
where $\Pi(\pi_{X},\pi_{Y})$ is the set of joint probability measures on $\mathcal{X}\times\mathcal{Y}$ with marginals equal to $\pi_{X},\pi_{Y}$ respectively. The existence of the minimizers are well-known; see, e.g., [64, Theorem 1.3]. The minimization problem in (3.4) is called the Monge–Kantorovich’s optimal transport problem [64]. If $\mathcal{X}=\mathcal{Y}$, and $d$ is a complete metric on $\mathcal{X}$ that induces the topology on this space (i.e., $(\mathcal{X},d)$ is a Polish metric space), and $c=d^{p}$ with $p\geq 1$, then $\mathsf{W}_{p}(\pi_{X},\pi_{Y}):=\mathsf{C}(\pi_{X},\pi_{Y})^{1/p}$ is the $p$-th Wasserstein metric between $\pi_{X}$ and $\pi_{Y}$.
For a probability measure $\pi_{W}$ and two transition probability $\pi_{X|W}$ and $\nu_{X|W}$, the conditional relative entropy of $\pi_{X|W}$ w.r.t. $\nu_{X|W}$ given $\pi_{W}$ is defined as
(3.5) $D(\pi_{X|W}\|\nu_{X|W}|\pi_{W}):=\int D(\pi_{X|W=w}\|\nu_{X|W=w}){\text{\rm d}}\pi_{W}(w).$
The conditional optimal transport cost between $\pi_{X|W}$ and $\pi_{Y|W}$ given $\pi_{W}$ is defined as
(3.6) $\mathsf{C}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\int\mathsf{C}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w).$
In particular, for $c=d^{p}$, we denote
(3.7) $\mathsf{W}_{p}(\pi_{X|W},\pi_{Y|W}|\pi_{W}):=\left(\int\mathsf{W}_{p}^{p}(\pi_{X|W=w},\pi_{Y|W=w}){\text{\rm d}}\pi_{W}(w)\right)^{1/p}.$
An important example is the following one.
Example 5 (Metric-Valued Cost).
$\mathcal{X}=\mathcal{Y}$ equipped with a metric $d$ is a Polish metric space, and the cost function is set to $c=d^{p}$ for $p\geq 1$, i.e., $\mathsf{C}=\mathsf{W}_{p}^{p}$.
To characterize the asymptotic concentration exponent, we need the following assumption. Define the $(\mathcal{X},\epsilon)$-smooth cost function w.r.t. $c$ as
$c_{\mathcal{X},\epsilon}(x,y):=\inf_{x^{\prime}:d(x,x^{\prime})\leq\epsilon}c(x^{\prime},y).$
By definition, $c_{\mathcal{X},\epsilon}(x,y)\leq c_{\mathcal{X},0}(x,y)=c(x,y)$, and by the lower semicontinuity of $c$, $\lim_{\epsilon\downarrow 0}c_{\mathcal{X},\epsilon}(x,y)\geq c(x,y)$. So, $\lim_{\epsilon\downarrow 0}c_{\mathcal{X},\epsilon}(x,y)=c(x,y)$ pointwise.
Assumption 3.
We assume that there is a function $\delta(\epsilon):(0,\infty)\to(0,\infty)$ vanishing as $\epsilon\downarrow 0$ such that
(3.8) $c_{\mathcal{X},\epsilon}(x,y)\geq(1-\delta(\epsilon))c(x,y)-\delta(\epsilon),\;\forall(x,y).$
Assumption 3 is satisfied by Example 5. We now verify this point.

### v4

@@ 3.1. Problem Formulation and Preliminaries
    For any $(\alpha,\tau)$ in the interior of $\mathrm{dom}\psi:=\left\{(\alpha,\tau):\psi(\alpha,\tau|\nu_{X},\nu_{Y})<\infty\right\}$, it holds that
    (3.7) $\liminf_{n\to\infty}E_{n}(\alpha,\tau|\nu_{X},\nu_{Y})\geq\psi(\alpha,\tau|\nu_{X},\nu_{Y}).$
    In particular, if further assume $\nu_{X}=\nu_{Y}$, then (3.7) holds for any $\alpha\geq 0,\tau\geq 0$.
Remark 6.
Theorem 7 still holds if the constraint $\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\})$ in (3.4) is replaced by $\pi_{XW}\in P_{\mathrm{c}}(\mathcal{X}\times\{0,1\})$.
Why is the asymptotic exponent determined by the relative entropy and the Wasserstein metric? The intuition is as follows. For a distribution $\pi$, denote $B_{\epsilon]}(\pi)$ as the closed ball of center $\nu$ and radius $\epsilon$ under the Lévy–Prokhorov metric. Let $\mathsf{L}_{n}:\mathbf{x}\in\mathcal{X}^{n}\mapsto\mathsf{L}_{\mathbf{x}}\in\mathscr{P}(\mathcal{X})$ be the empirical measure map, where $\mathsf{L}_{\mathbf{x}}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ is the empirical measure of $x$, and $\mathscr{P}(\mathcal{X})$ is the set of probability measures on $\mathcal{X}$. For a Polish space $\mathcal{X}$ and $\epsilon>0$, the empirically $\epsilon$-typical set of $\pi$ [53] is defined as $\mathcal{T}_{\epsilon}^{(n)}(\pi):=\mathsf{L}_{n}^{-1}(B_{\epsilon]}(\pi)).$
For an empirically $\epsilon$-typical set $A=\mathsf{L}_{n}^{-1}(B_{\epsilon}(\pi_{X}))$ (i.e., $A=\{\mathbf{x}\in\mathcal{X}^{n}:\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$), roughly speaking, by Sanov’s theorem it holds that $\nu_{X}^{\otimes n}(A)\approx e^{-nD(\pi_{X}\|\nu_{X})}$ for $\epsilon$ chosen sufficiently small, which is no smaller than $e^{-n\alpha}$ if $D(\pi_{X}\|\nu_{X})\leq\alpha$. The $\sqrt{n\tau}$-enlargement of this set can be written as
$A^{\sqrt{n\tau}}$ $=\{\mathbf{y}:\frac{1}{n}d_{n}^{2}(\mathbf{x},\mathbf{y})\leq\tau,\exists\mathbf{x}\in A\}$
$=\{\mathbf{y}:\mathbb{E}_{\mathsf{L}_{\mathbf{x},\mathbf{y}}}[d^{2}(X,Y)]\leq\tau,\mathsf{L}_{\mathbf{x}}\approx\pi_{X},\exists\mathbf{x}\}$
(3.8) $=\{\mathbf{y}:\mathsf{W}_{2}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})\leq\sqrt{\tau},\mathsf{L}_{\mathbf{x}}\approx\pi_{X}\}$
$\approx\mathsf{L}^{-1}\{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}\},$
where (3.8) is due to Birkhoff’s theorem, $\min_{\sigma}\mathbb{E}_{\mathsf{L}_{\sigma(\mathbf{x}),\mathbf{y}}}[d^{2}(X,Y)]=\mathsf{W}_{2}^{2}(\mathsf{L}_{\mathbf{x}},\mathsf{L}_{\mathbf{y}})$, with $\sigma$ denoting a permutation operator [72, Page 5]. By Sanov’s theorem again,
$\nu_{Y}^{\otimes n}(A^{\sqrt{n\tau}})\approx\exp\{-n\inf_{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}}D(\pi_{Y}\|\nu_{Y})\}.$
Optimizing over $\pi_{X}$ such that $D(\pi_{X}\|\nu_{X})\leq\alpha$, we obtain the isoperimetric exponent for $A$,
$\sup_{\pi_{X}:D(\pi_{X}\|\nu_{X})\leq\alpha}\inf_{\pi_{Y}:\mathsf{W}_{2}(\pi_{X},\pi_{Y})\leq\sqrt{\tau}}D(\pi_{Y}\|\nu_{Y}).$
By product construction of such $A$’s (forming a conditional empirically typical set), we obtain the exponent $\psi(\alpha,\tau|\nu_{X},\nu_{Y})$.
The intuition above provides a proof idea for Statement 2 of Theorem 7. However, Statement 1 is highly nontrivial and its proof relies on the so-called the inherently typical subset lemma [2]. This lemma enables us to tailor a set into a new one that is “regular” in some sense, so that we can analyze the asymptotic expression of the boundary size of the set.
3.2. Riemannian Energy Measure Spaces Satisfying $\text{BE}(K,\infty)$ and $\operatorname{RCD}(K,\infty)$-spaces
Our goal in the whole section is to generalize Statement 1 of Theorem 7 to probability measures satisfying Condition 2. However, since the class of Polish metric measure spaces is too broad to allow for straightforward derivation of bounds, we restrict our analysis to a subclass of Polish spaces with favorable properties—namely, Riemannian energy measure spaces satisfying the $\text{BE}(K,\infty)$ condition, or equivalently, $\text{RCD}(0,\infty)$-spaces. This choice enables us to use the well-defined heat flow to establish bounds for variant isoperimetry. To this end, we now introduce these two concepts; for further details, we refer the reader to [7, 8].
We consider an Energy measure space $(\mathcal{X},d,\nu)$, where $(\mathcal{X},d)$ is a Polish metric space, $\nu$ is a Borel probability measure with full support, and the underlying $\sigma$-algebra is the $\nu$-completion of the Borel $\sigma$-algebra. The structure is governed by the Cheeger energy $\text{Ch}:L^{2}(\mathcal{X},\nu)\to[0,\infty]$ given in (2.2), defined as the $L^{2}$-relaxation of the local Lipschitz constant. The space is assumed to be infinitesimally Hilbertian in the following sense.
  • $\mathcal{E}=2\text{Ch}$ is a quadratic form that constitutes a strongly local, symmetric Dirichlet form $(\mathcal{E},\mathbb{V})$ with domain $\mathbb{V}:=D(\mathcal{E})=\bigl\{f\in L^{2}(\mathcal{X},\nu):\mathcal{E}(f)<\infty\bigr\}$.
Define the set $\mathbb{G}\subseteq\mathbb{V}$ such that for $f\in\mathbb{G}$, the linear form
$\boldsymbol{\Gamma}[f;\varphi]:=\mathcal{E}(f,f\varphi)-\tfrac{1}{2}\mathcal{E}(f^{2},\varphi),\;\varphi\in\mathbb{V}_{\infty}:=\mathbb{V}\cap L^{\infty}(\mathcal{X},\nu)$
is represented by the $L_{+}^{1}$ density $\Gamma(f)=|\nabla f|_{*}^{2}$, satisfying $\mathcal{E}(f,f)=\int\Gamma(f)\,{\text{\rm d}}\nu$. Define $\Gamma(f,g)=\frac{1}{4}\Gamma(f+g)-\frac{1}{4}\Gamma(f-g)$ which is known as the carré du champ operator. By the uniform continuity property, we extend $\boldsymbol{\Gamma}$ to a real multilinear map defined in $\mathbb{V}\times\mathbb{V}\times\mathbb{V}_{\infty}$, which satisfies
$\mathbf{\Gamma}[f,g;\varphi]=\int\Gamma(f,g)\varphi\,{\text{\rm d}}\nu\quad\text{if }f,g\in\mathbb{G}.$
Define $\mathcal{L}:=\{\psi\in\mathbb{G}:\Gamma(\psi)\leq 1\quad\nu\text{-a.e. in }\mathcal{X}\}$ and $\mathcal{L}_{C}:=\mathcal{L}\cap C(\mathcal{X})$. The compatibility between the metric $d$ and the energy $\mathcal{E}$ is enforced by the following structural assumptions:
  • Identification: $\mathcal{L}=\mathcal{L}_{C}$. That is, every function with a gradient density bounded by one $\nu$-a.e. admits a continuous representative.
  • Lipschitz Consistency: Every function in $\mathcal{L}_{C}$ is $1$-Lipschitz with respect to $d$.
The above three assumptions turns the space $(\mathcal{X},d,\nu)$ described above into a Riemannian Energy measure space in which the intrinsic distance $d_{\mathcal{E}}$ induced by the Dirichlet form coincides with the distance $d$, generalizing smooth Riemannian manifolds to the possibly non-smooth setting.
The Dirichlet form $\mathcal{E}$ induces a densely defined self-adjoint operator $\Delta_{\mathcal{E}}:D(\Delta_{\mathcal{E}})\subset\mathbb{V}\to L^{2}(\mathcal{X},\nu)$ defined by the integration by parts formula $\mathcal{E}(f,g)=-\int g\Delta_{\mathcal{E}}f\,{\text{\rm d}}\nu$ for all $g\in\mathbb{V}$. This self-adjoint operator generates the heat flow $(P_{t})_{t\geq 0}$, an analytic Markov semigroup characterized by $\frac{\mathrm{d}}{\mathrm{d}t}P_{t}f=\Delta_{\mathcal{E}}P_{t}f$, a specialization of the heat equation in (2.4) in the Riemannian Energy measure space.
Bakry–Émery conditions are formulated via the iterated gradient form $\boldsymbol{\Gamma}_{2}$. For $(f,\varphi)\in D(\boldsymbol{\Gamma}_{2})$, define:
$\boldsymbol{\Gamma}_{2}[f;\varphi]:=\tfrac{1}{2}\boldsymbol{\Gamma}[f;\Delta_{\mathcal{E}}\varphi]-\boldsymbol{\Gamma}[f,\Delta_{\mathcal{E}}f;\varphi],$
where
$D(\Gamma_{2}):=\left\{(f,\varphi)\in D(\Delta_{\mathcal{E}})\times D(\Delta_{\mathcal{E}}):\Delta_{\mathcal{E}}f\in\mathbb{V},\,\varphi,\Delta_{\mathcal{E}}\varphi\in L^{\infty}(\mathcal{X},\nu)\right\}.$
We can write
$\mathbf{\Gamma}_{2}[f;\varphi]=\int\left(\frac{1}{2}\Gamma(f)\Delta_{\mathcal{E}}\varphi-\Gamma(f,\Delta_{\mathcal{E}}f)\varphi\right){\text{\rm d}}\nu\quad\text{if }(f,\varphi)\in D(\mathbf{\Gamma}_{2}),\ f,\Delta_{\mathcal{E}}f\in\mathbb{G}.$
The space satisfies the Bakry–Émery condition $\text{BE}(K,\infty)$ if for every $\varphi\in D(\Delta_{\mathcal{E}})$ with $\varphi\geq 0$:
$\boldsymbol{\Gamma}_{2}[f;\varphi]\geq K\boldsymbol{\Gamma}[f;\varphi].$
In the setting of smooth weighted Riemannian manifolds, this condition can be written as $\frac{1}{2}\Gamma(f)\Delta_{\mathcal{E}}-\Gamma(f,\Delta_{\mathcal{E}}f)\geq K\Gamma(f)$, or equivalently, $\mathrm{Ric}_{V}\geq Kg$ for probability measure $\nu=e^{-V}\mathrm{vol}$. Under $\text{BE}(K,\infty)$, it is known that $\mathbb{G}=\mathbb{V}$.
The Riemannian Energy measure space $(\mathcal{X},d,\nu)$ (with $d=d_{\mathcal{E}}$) satisfies $\text{BE}(K,\infty)$ if and only if it is an $\text{RCD}(K,\infty)$ space [8, Theorem 4.17]. From the perspective of the geometry of optimal transport, this equivalence implies that the relative entropy functional $D(\cdot\|\nu)$ is not merely $K$-geodesically convex on the Wasserstein space $(\mathscr{P}_{2}(\mathcal{X}),\mathsf{W}_{2})$, but that this convexity is robust: it extends to hold along all interpolated measures induced by weighted optimal geodesic plans (or test plans).
A fundamental consequence of the $\operatorname{RCD}(K,\infty)$ property (equivalently, the Riemannian Energy measure space with $\text{BE}(K,\infty)$ property) is the identification of the metric $\mathsf{W}_{2}$-gradient flow $\mathcal{H}_{t}$ with the analytic $L^{2}$-heat flow $P_{t}$. For any $f\in L^{2}(\mathcal{X},\nu)$ such that $f\nu\in\mathscr{P}_{2}(\mathcal{X})$, the identification $\mathcal{H}_{t}(f\nu)=(P_{t}f)\nu$ holds. Furthermore, the heat flow admits the representation:
$(P_{t}f)\nu=\int f(x)\mathcal{H}_{t}(\delta_{x})\,{\text{\rm d}}\nu(x),$
characterizing the evolved measure as a weighted superposition of the evolved Dirac masses.
3.3. Variant Isoperimetry for Noncompactly-Supported Measures
We next extend Statement 1 of Theorem 7 to the Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ and Condition 2 (equivalently, a $\mathrm{RCD}(0,\infty)$-space satisfying Condition 2). To this end, we need the following conditions on the covering bound, volume ratio bound, and curvature-dimension bound.
Condition 4 (Covering Bound).
Let $x_{0}\in\mathcal{X}$ be a fixed point. Then, for any $n\geq 1$, $\lambda\geq 1$, and $r>0$, there is cover of the large ball $B_{\sqrt{n}\lambda r}(\mathbf{x}_{0})$ with $\mathbf{x}_{0}=(x_{0},...,x_{0})\in\mathcal{X}^{n}$ that consists of small balls $B_{\sqrt{n}r}(\mathbf{x}_{i})$ with $\mathbf{x}_{i}\in\mathcal{X}^{n}$ and $1\leq i\leq N$, i.e., $B_{\sqrt{n}\lambda r}(\mathbf{x}_{0})\subseteq\bigcup_{1\leq i\leq N}B_{\sqrt{n}r}(\mathbf{x}_{i})$, where $N=N_{n,\lambda,r}$ satisfies
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n,\lambda,r}=0,\;\forall\lambda\geq 1.$
Condition 5 (Volume Ratio Bound).
There is a nonnegative (not necessarily finite or probability) measure $\mathfrak{m}$ on $\mathcal{X}^{n}$ such that for all $\lambda\geq 1$,
(3.9) $\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N_{n}(\sqrt{n}\lambda r,\sqrt{n}r)=0,$
where
(3.10) $N_{n}(R,r)=\sup_{\mathbf{x}\in\mathcal{X}^{n}}\frac{\mathfrak{m}(B_{R}(\mathbf{x}))}{\mathfrak{m}(B_{r}(\mathbf{x}))}.$
Condition 6 (Curvature-Dimension Bound).
Suppose $\mathcal{X}$ is a subset of a Riemannian manifold $(\mathcal{M},g)$ or the interior of $\mathcal{X}$ is a Riemannian manifold $(\mathcal{M},g)$. There is a nonnegative (not necessarily finite or probability) measure $\mathfrak{m}=e^{-U}\mathrm{vol}$ on manifold $\mathcal{M}$ with $U\in C^{2}(\mathcal{M})$ satisfying $\mathrm{CD}(K,N)$ for some $N\in(k,\infty)$ and $K\in\mathbb{R}$ with $k=\mathrm{dim}(\mathcal{M})$:
$\mathrm{Ric}_{U}^{N}:=\mathrm{Ric}_{g}+\mathrm{Hess}_{g}U-\frac{\nabla U\otimes\nabla U}{N-k}\geq Kg.$
The conditions specified above and Condition 3 satisfy the following implication relation.
Proposition 3.
It holds that Condition 5 $\Longrightarrow$ Condition 4. Moreover, if $\mathcal{X}$ is a subset of a Riemannian manifold $(\mathcal{M},g)$ or the interior of $\mathcal{X}$ is a Riemannian manifold $(\mathcal{M},g)$, then Condition 3 $\Longrightarrow$ Condition 6 $\Longrightarrow$ Condition 5 $\Longrightarrow$ Condition 4.
Define
$\Upsilon_{\alpha}(s)=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Define
(3.11) $\bar{\psi}(\alpha,\tau)$ $=\begin{cases}\Upsilon_{\alpha}^{-1}(\sqrt{\tau}),&\Upsilon_{\alpha}(0)\geq\sqrt{\tau}\\ 0,&\Upsilon_{\alpha}(0)<\sqrt{\tau}\end{cases},$
where for each $\alpha$, denote $\Upsilon_{\alpha}^{-1}$ as the inverse of $\Upsilon_{\alpha}$. We are ready to extend Statement 1 of Theorem 7 to all probability measures satisfying Condition 2 or equivalently admitting a standard log-Sobolev inequality.
Theorem 8 (Asymptotic Exponent for Noncompactly-Supported Measures).
Let $(\mathcal{X},d,\nu)$ be a Riemannian Energy measure space satisfying $\text{BE}(K,\infty)$ (equivalently, it is a $\mathrm{RCD}(0,\infty)$-space). Suppose that this space satisfies Condition 2 and it admits a $\theta$-nonlinear log-Sobolev inequality. If the condition in (1.10) holds, we further assume Condition 4 (or Conditions 3, 6, or 5) holds. Then, for any $\alpha\geq 0,\tau\geq 0$,
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\bar{\psi}(\alpha,\tau).$
Remark 7.
Under $\mathrm{RCD}(0,\infty)$ (equivalently, the Riemannian Energy setting with $\mathrm{BE}(0,\infty)$) for a probability measure $\nu\in\mathscr{P}_{2}(\mathcal{X})$, Condition 2 holds if and only if it admits a standard log-Sobolev inequality [9, Theorem 3.2].
Remark 8.
Define
$\psi(\alpha,\tau)$ $:=\psi(\alpha,\tau|\nu,\nu)$
$=\sup_{\pi_{XW}\in\mathscr{P}(\mathcal{X}\times\{0,1\}):D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}$
(3.12) $\qquad\inf_{\pi_{Y|W}:\mathsf{W}_{2}(\pi_{X|W},\pi_{Y|W}|\pi_{W})\leq\sqrt{\tau}}D(\pi_{Y|W}\|\nu|\pi_{W}),$
and
$\tilde{\psi}(\alpha,\tau)=\frac{1}{2}\left[\sqrt{2\alpha}-\sqrt{K_{LS}^{+}\tau}\right]_{+}^{2},$
where recall the definition of $K_{LS}^{+}$ in (1.15). By checking our proof, the following bound can be obtained:
$\limsup_{n\to\infty}E_{n}(\alpha,\tau)\leq\sup_{\alpha_{0}+\alpha_{1}\leq\alpha}\inf_{\tau_{0}+\tau_{1}\leq\tau}\psi(\alpha_{0},\tau_{0})+\tilde{\psi}(\alpha_{1},\tau_{1}).$
This bound is more consistent with the lower bound in Statement 2 of Theorem 7, and it coincides with the latter if $\psi\leq\tilde{\psi}$.
3.4. Proof of Proposition 3
By choosing $\mathfrak{m}$ as the Riemannian volume measure, Condition 3 $\Longrightarrow$ Condition 6 is obvious. We next show that Condition 5 $\Longrightarrow$ Condition 4.
For a subset $A\subseteq\mathcal{M}^{n}$, denote by $N(A,r)$ the minimal covering number (i.e., the smallest number of balls of radius $r$ needed to cover $A$) and by $M(A,r)$ the maximal packing number (i.e., the largest number of disjoint balls of radius $r$ whose centers lie in $A$). By the standard lemma for metric spaces, $N(A,r)\leq M(A,r/2).$ Setting $A=B_{R}(\mathbf{x})$ yields that $N(B_{R}(\mathbf{x}),r)\leq M(B_{R}(\mathbf{x}),r/2).$
For any ball $B(\mathbf{x}_{i},r/2)$ in the packing for $M(B_{R}(\mathbf{x}),r/2)$, since $p_{i}\in B_{R}(\mathbf{x})$, all such balls are contained in $B_{R+r/2}(\mathbf{x})$. Let $i^{*}$ minimize $\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i}))$ over all $i$. Combining this with the fact that $B_{R+r/2}(\mathbf{x})\subseteq B_{2R+r/2}(\mathbf{x}_{i^{*}})$, we obtain:
$M(B_{R}(\mathbf{x}),r/2)\cdot\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i^{*}}))\leq\mathfrak{m}^{\otimes n}(B_{R+r/2}(\mathbf{x}))\leq\mathfrak{m}^{\otimes n}(B_{2R+r/2}(\mathbf{x}_{i^{*}})).$
Therefore, for any $\mathbf{x}$ and any $R>r>0$,
$N(B_{R}(\mathbf{x}),r)\leq M(B_{R}(\mathbf{x}),r/2)\leq\frac{\mathfrak{m}^{\otimes n}(B_{2R+r/2}(\mathbf{x}_{i^{*}}))}{\mathfrak{m}^{\otimes n}(B_{r/2}(\mathbf{x}_{i^{*}}))}\leq N_{n}(2R+r/2,r/2).$
Substituting $(R,r)\leftarrow(\sqrt{n}\lambda r,\sqrt{n}r)$ into the inequality above and by the condition (3.9), we obtain that
$\lim_{r\to\infty}\limsup_{n\to\infty}\frac{1}{nr^{2}}\ln N(B_{\sqrt{n}\lambda r}(\mathbf{x}_{0}),\sqrt{n}r)=0,\;\forall\lambda\geq 1.$
Therefore, Condition 4 holds.
We now prove that Condition 6 $\Longrightarrow$ Condition 5. Let $\mathfrak{m}=e^{-U}\mathrm{vol}$ be the nonnegative measure on manifold $\mathcal{M}$ given in Condition 6. Then, $\mathfrak{m}^{\otimes n}=e^{-U^{\oplus n}}\mathrm{vol}_{n}$ is a nonnegative measure on manifold $\mathcal{M}^{n}$ with $U^{\oplus n}\in C^{2}(\mathcal{M}^{n})$ such that $\mathrm{Ric}_{U^{\oplus n}}^{nN}\geq nK$.
For a point $p$ in $\mathcal{M}$ and $\xi\in T_{p}\mathcal{M}$ such that $|\xi|=1$, let $\gamma_{\xi}(t)=\exp_{p}(t\xi)$ be the geodesic starting at $p$ with direction $\xi$, and let $\lambda_{p}(t,\xi)=\lambda_{p}^{U}(t,\xi)$ be the density of the weighted measure $\mathfrak{m}$ at the point $\gamma_{\xi}(t)$. The density $\lambda_{p}(t,\xi)$ is defined to be zero whenever $t$ exceeds the distance to the cut point in the direction $\xi$. This density incorporates both the Riemannian volume distortion (Jacobian of the exponential map) and the weight function $e^{-U}$. Let $\lambda^{(K)}(t)$ be the Riemannian volume density in the $(\mathrm{dim}(\mathcal{M})+N)$-dimensional model space $\mathcal{M}^{(K)}$ with constant Ricci curvature $K$. Following [60, Theorem 7], under the condition $\mathrm{Ric}_{U}^{N}\geq Kg$, it holds that for any $\xi$, $\frac{\lambda_{p}(t,\xi)}{\lambda^{(K)}(t)}$ is non-increasing in $t$ for $\gamma_{\xi}(t)\notin\mathrm{cut}(p)$, where $\mathrm{cut}(p)$ is the cut locus of the Riemannian manifold $M$ with respect to the point $p$.
Let $A(p,t)$ be the $(\mathrm{dim}(\mathcal{M})-1)$-dimensional $\mathfrak{m}$-weighted area of a geodesic sphere of radius $t$ at center $p$ in $M$. Then, for any $t>0$,
$A(p,t)$ $=\int_{S_{p}\mathcal{M}}\lambda_{p}(t,\xi)\,{\text{\rm d}}\mathrm{area}(\xi)$
where $S_{p}\mathcal{M}$ is the unit sphere in the tangent space $T_{p}\mathcal{M}$ at point $p$, and ${\text{\rm d}}\mathrm{area}(\xi)$ is the standard area element of the unit (n1)(n-1)-sphere in the tangent space.
Let $A^{(K)}(t)$ be the $(\mathrm{dim}(\mathcal{M})-1)$-dimensional Riemannian area of a geodesic sphere of radius $t$ at center $p$ in the $\mathrm{dim}(\mathcal{M})$-dimensional model space $\mathcal{M}^{(K)}$ with constant Ricci curvature $K$. Then,
$A^{(K)}(t)$ $=\lambda^{(K)}(t)\cdot|S_{p}\mathcal{M}^{(K)}|,$
where $S_{p}\mathcal{M}^{(K)}$ is the unit sphere in the tangent space $T_{p}\mathcal{M}^{(K)}$ at point $p$, and $|S_{p}\mathcal{M}^{(K)}|$ is the standard area of the unit (n1)(n-1)-sphere in the tangent space. Note that the value of $|S_{p}\mathcal{M}^{(K)}|$ is independent of $p$.
Observe that
$\frac{A(p,t)}{A^{(K)}(t)}=\frac{1}{|S_{p}\mathcal{M}^{(K)}|}\int_{S_{p}\mathcal{M}}\frac{\lambda_{p}(t,\xi)}{\lambda^{(K)}(t)}\,{\text{\rm d}}\mathrm{area}(\xi),$
which is non-increasing in $t$. The volume of the ball with center $\mathbf{p}=(p_{1},...,p_{n})$ and radius $R$ in the space $\mathcal{M}^{n}$ is
$V_{n}(\mathbf{p},R)=\int_{\sum_{i=1}^{n}\rho_{i}^{2}\leq R^{2}}\prod_{i=1}^{n}A(p_{i},\rho_{i}){\text{\rm d}}\rho_{i}.$
Similarly, the volume of the ball with radius $R$ (at any center) in the space $(\mathcal{M}^{(K)})^{n}$ is
$V_{n}^{(K)}(R)$ $=\int_{\sum_{i=1}^{n}\rho_{i}^{2}\leq R^{2}}\prod_{i=1}^{n}A^{(K)}(\rho_{i}){\text{\rm d}}\rho_{i}.$
We have the following key lemma.
Lemma 1.
Given any $\mathbf{p}$, $f(R)=\frac{V_{n}(\mathbf{p},R)}{V_{n}^{(K)}(R)}$ is non-increasing.